OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: x – 5y > –7 3x + 4y less than or equal to –8 (4, 5) (–2, 4) (–1, –1) (3, –8)

OpenStudy (jdoe0001):

try the choices out :), plug them in, see if they give a valid inequality let's try the first one (4, 5) $$\bf x-5y > -7\\ 3x+4y \le -8\\ --------------\\ (4)-5(5) > -7\\ 3(4)+4(5) \le -8\\ -21 > -7 \text{ # false -21 is really smaller}\\ 32 \le -8 \ \ \ \ \ \text{ # false, 32 is really bigger} $$

OpenStudy (anonymous):

so it's not A

OpenStudy (jdoe0001):

right

OpenStudy (anonymous):

is the answer B?

OpenStudy (jdoe0001):

let's see (-2, 4) (-2) - 5(4) > -7 -22 > -7 # no dice

OpenStudy (jdoe0001):

-22 is really smaller than -7

OpenStudy (anonymous):

ohhh ok ok

OpenStudy (anonymous):

so it's D then... (i think you can see the pattern i'm doing here)

OpenStudy (jdoe0001):

(3, -8) (3) - 5(-8) > -7 43 > -7 # indeed 3(3) + 4(-8) <= -8 -23 <= -8 # yes it's

OpenStudy (anonymous):

yayyy